N2H4 (l) + CH4O (l) CH2O (g) + N2 (g) + 3H2 (g)


1
Expert's answer
2021-11-08T08:23:57-0500


HESS' LAW


N2H4(l) + H2(g) ==> 2NH3(g) ... TARGET EQUATION


Given:

(1) N2H4(l) + CH4O(l) ==> CH2O(g) + N2(g) + 3H2(g) ... ∆H = -37 kJ

(2) N2(g) + 3H2(g) ==> 2NH3(g) ... ∆H = -46 KJ

(3) CH4O(l) ==> CH2O(g) + H2(g) ... ∆H = -65 kJ


Procedure:

Copy (1): N2H4(l) + CH4O(l) ==> CH2O(g) + N2(g) + 3H2(g) ... ∆H = -37 kJ

Copy (2): N2(g) + 3H2(g) ==> 2NH3(g) ... ∆H = -46 KJ

Reverse (3): CH2O(g) + H2(g) => CH4O(l) ... ∆H = +65 kJ (note to reverse sign)

Add them up and combine and/or cancel like terms to get...

N2H4(l) + CH4O(l) + N2(g) + 3H2(g) + CH2O(g) + H2(g) => CH2O(g) + N2(g) + 3H2(g) + 2NH3(g) + CH4O(l)

N2H4(l) + H2(g) ==> 2NH3(g) ... TARGET EQUATION

∆H = -37 kJ + -46 kJ + 65 kJ

∆H = -18 kJ